### v1

# Mut4All: Fuzzing Compilers via LLM-Synthesized Mutators Learned from Bug Reports

Abstract.
Mutation-based fuzzing has proven effective in uncovering compiler bugs. However, designing high-quality mutation operators (i.e., mutators) for modern programming languages that feature complex language constructs (e.g., templates and macros) remains challenging. Existing methods rely heavily on manual effort, either to craft a limited set of simple mutators or to inspect and correct automatically generated ones. This reliance limits the expressiveness of mutators, hampers scalable mutator synthesis, and consequently restricts cross-language generalizability as mutators are often tailored to specific languages.
In this work, we present Mut4All, a fully automated and language-agnostic framework that synthesizes mutation operators by leveraging Large Language Models (LLMs) and compiler-specific knowledge extracted from bug reports on open-source repositories. It is a multi-agent framework comprising three specialized LLM agents: (1) a mutator invention agent, which identifies mutation targets and generates mutator names, signatures, and descriptions based on compiler-related insights (e.g., language features explicitly mentioned in bug reports); (2) a mutator implementation synthesis agent, which is enhanced via fine-tuning to generate initial mutator implementations; and (3) a mutator refinement agent, which verifies and corrects the synthesized mutators using feedback from unit tests. This pipeline analyzes 1000 bug reports (500 each for Rust and C++), and synthesizes 319 and 403 mutators for them. Each mutator only costs approximately \$0.08 using GPT-4o APIs. Using these mutators, our customized fuzzer has reported 62 bugs in the Rust compilers (rustc and gccrs), 38 of which were confirmed as previously unknown, with 7 already fixed by developers. It has also reported 34 bugs in C++ compilers (GCC and Clang), including 16 confirmed as new ones, 1 of which has been resolved. Experimental results also show that Mut4All discovers significantly more unique crashes than existing approaches, and achieves the highest coverage on Rust and the second-highest on C++.
Compiler, Mutation-Based Compiler Fuzzing, Compiler Bugs, Rust, C++
copyright: noneccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Compilers
1. Introduction
Compiler fuzzing (Chen et al., 2020; Marcozzi et al., 2019; Chen et al., 2016a) has proven effective in improving the reliability of production compilers such as GCC and rustc, uncovering numerous bugs over the past decades. Existing methods can be broadly categorized into two types: generation-based and mutation-based. Generation-based fuzzers, such as Csmith (Yang et al., 2011) and YARPGen (Livinskii et al., 2020, 2023), generate syntactically valid programs from scratch using the compiler’s input grammar. However, ensuring the semantic correctness of these test programs often requires substantial manual effort to design templates, particularly for modern programming languages with intricate features such as ownership and macros, which limits their scalability and adaptability.
In contrast, mutation-based fuzzers such as GrayC (Even-Mendoza et al., 2023) and MetaMut (Ou et al., 2024), which generate new test programs by applying mutation operators (i.e., mutators) to existing real-world programs (i.e., seed programs), are often preferred for compiler fuzzing, as they leverage existing code to avoid the need for designing complex templates and preserve both the syntactic and semantic correctness of test programs. That being said, their effectiveness heavily depends on the quality and expressiveness of the mutators they employ. Moreover, existing mutation-based fuzzers still face several major challenges in mutator design and implementation:
C1: Limited Expressiveness. Most existing mutation-based fuzzers define mutators as compositions of simple operators, such as arithmetic adjustments (e.g., replacing + with -) or control-flow rewrites (e.g., converting switch statements into if-else chains), as exemplified by GrayC (Even-Mendoza et al., 2023) and, to a more advanced extent, MetaMut (Ou et al., 2024). While effective at exposing bugs in basic language constructs, these mutators inherently lack expressiveness—the ability to construct program variants that incorporate advanced language features and engage deeper layers of compiler logic, such as trait objects in Rust or templates in C++, both of which have received limited attention in prior studies (Yang et al., 2024b; Even-Mendoza et al., 2023; Ou et al., 2024; Wang and Jung, 2024). This limitation often results in missed opportunities to uncover subtle bugs that arise from intricate interactions between language features, which are prevalent in realistic programs (Wang et al., 2025).
C2: Significant Manual Effort. Existing mutation-based compiler fuzzers are mostly semi-automatic and still require active involvement from compiler experts. Tools like GrayC (Even-Mendoza et al., 2023) and DIE (Park et al., 2020) rely on manually crafted mutators, incurring high engineering costs—the former implements 13 mutators using over 8,000 lines of C++, while the latter requires more than 16,000 lines for its mutation engine. More recent approaches leverage Large Language Models (LLMs), which have demonstrated strong capabilities across programming tasks (Hou et al., 2024; Wang et al., 2024; Chen et al., 2025a), for mutator generation. MetaMut (Ou et al., 2024) guides the LLM to assemble mutators from around 100 predefined AST-level operations, while Clozemaster (Gao et al., 2025) fine-tunes LLMs to complete masked syntax elements. While this significantly reduces manual implementation, it still involves non-trivial effort in prompt engineering, as well as validating and refining the generated code. For instance, MetaMut reports over two weeks of manual effort to repair and refine just 68 mutators (Ou et al., 2024). Although promising, these methods still have room for improvement in scalability and automation, especially when scaling to support a large number of mutators for diverse compiler features.
C3: Limited Cross-Language Generalizability. Ideally, a compiler fuzzer should work across multiple programming languages and compilers for broader applicability (Xia et al., 2024). That said, due to substantial differences in language design and compiler infrastructures, mutators’ implementation is often tightly coupled to a specific compiler. Even when designed to be generic, mutators that work well for one compiler may not be effective—or even functional—for another, making cross-language reuse difficult. Most existing fuzzers are designed for a single language, with mutators tailored to specific compiler infrastructures (e.g., GrayC (Even-Mendoza et al., 2023) and Clozemaster (Gao et al., 2025) only work for C and Rust, respectively). Extending them to new targets is inherently challenging, often requiring thousands of lines of code to accommodate differences in syntax, type systems, and compiler internals. While the rise of LLMs trained on multilingual code offers promise for cross-language transfer (Xia et al., 2024), current mutation-based fuzzers (even MetaMut and Clozemaster) have yet to fully exploit this capability. Achieving high-quality mutator transfer and reuse across languages remains an open challenge.
Our work.  We introduce Mut4All, a fully automated and language-agnostic framework for synthesizing mutation operators to support mutation-based compiler fuzzing. The core idea behind Mut4All is to automate the entire mutator lifecycle, from discovery and design to implementation, by coordinating specialized LLM agents that extract compiler-specific insights from historical bug reports in open-source repositories (e.g., GitHub). These real-world bug reports help identify bug-prone language features and provide concrete test programs that reveal compiler failures, allowing LLMs to focus on patterns that have previously caused actual bugs and produce more targeted and effective mutations. At a high level, Mut4All (1) analyzes real-world bug reports to identify language features prone to compiler bugs and formulates corresponding mutator specifications, (2) synthesizes concrete mutator implementations based on these specifications, and (3) validates and refines the generated mutators to ensure their correctness and effectiveness.
Concretely, given a target compiler, the workflow begins by automatically crawling related bug reports from open-source repositories, which serve as the knowledge base for three LLM agents in Mut4All. These agents operate in a multi-agent fashion as follows:
  • Mutator Invention Agent: Analyzes historical bug reports to identify error-prone language features, such as those frequently associated with compiler bugs, and proposes mutator designs as specifications describing the intended transformations.
  • Mutator Implementation Synthesis Agent: Synthesizes concrete mutator code guided by the specifications and high-quality AST modification examples. It is also fine-tuned with a curated set of high-quality seed mutators to produce more accurate and diverse transformations.
  • Mutator Refinement Agent: Validates each synthesized mutator by applying it to bug-triggering programs collected from historical bug reports. If compilation or runtime errors occur, the agent automatically repairs it using error feedback. This process iterates until validation succeeds or a retry limit is reached.
Following this workflow, we apply Mut4All to two prominent system-level programming languages, Rust and C++, and automatically analyze 1,000 bug reports (500 each for Rust and C++) from two open-source repositories: rustc for Rust and Clang for C++. In total, Mut4All synthesizes 319 mutators for Rust and 403 mutators for C++ without any human intervention, at an average cost of approximately \$0.08 per mutator using GPT-4o APIs. These operators target complex and error-prone language features, such as trait object layouts [^1] https://doc.rust-lang.org/book/ch10-02-traits.html in Rust and templates [^2] https://en.cppreference.com/w/cpp/language/templates.html in C++, which remain underexplored in prior work (Even-Mendoza et al., 2023; Ou et al., 2024; Sharma et al., 2023; Wang and Jung, 2024; Yang et al., 2024b).
The synthesized mutators are then integrated into a customized fuzzer targeting production-level compilers, including rustc, gccrs, GCC, and Clang. Seeded with bug-triggering programs from historical reports and official test cases, the fuzzer uses crash and hang oracles, as well as differential testing across compilers of the same language (e.g., GCC and Clang for C++). We benchmark Mut4All against a range of state-of-the-art compiler fuzzers, including traditional fuzzers such as RustSmith (Sharma et al., 2023) and Rustlantis (Wang and Jung, 2024) for Rust, and CSmith (Yang et al., 2011), YARPGen (Livinskii et al., 2020; Li et al., 2023), and GrayC (Even-Mendoza et al., 2023) for C++. Additionally, we compare against recent LLM-based fuzzers, including Clozemaster (Gao et al., 2025) for Rust, as well as Fuzz4All (Xia et al., 2024) and MetaMut (Ou et al., 2024) for C++. The results indicate that Mut4All outperforms existing approaches by discovering more unique crashes, achieving the highest line coverage on Rust, and ranking second in coverage on C++. Particularly, compared with the approaches that directly leverage LLMs to mutate code (i.e., Clozemaster (Gao et al., 2025) and Fuzz4All (Xia et al., 2024)), our approach consistently performs better. In a 48-hour bug-hunting campaign, Mut4All uncovered a total of 96 bugs across both languages. For the Rust compilers (rustc and gccrs), Mut4All reported 62 bugs, of which 38 were confirmed as previously unknown and 7 had already been fixed by developers. For the C++ compilers (GCC and Clang), it reported 34 bugs, including 16 previously unknown and 1 already fixed. These results demonstrate that Mut4All can effectively synthesize high-quality mutators for complex language features, enabling practical bug discovery in real-world compilers.
Contributions.  Mut4All addresses the key limitations of existing mutation-based compiler fuzzers. To overcome the limited scope and expressiveness of existing mutators (C1), it automatically invents new mutators guided by compiler-specific insights from open-source bug reports, often covering advanced language features. Unlike prior approaches that require manual effort to design mutators (C2) and are typically restricted to a single language or compiler (C3), Mut4All leverages LLMs as the generation engine—minimizing human intervention and enabling cross-language transfer through LLMs’ multilingual knowledge. In summary, this paper makes the following contributions:
  • We propose Mut4All, the first fully automated multi-agent framework that synthesizes high-quality mutators for production-level compilers, guided by insights from historical compiler bug reports widely spread on open-source communities.
  • Mut4All synthesizes 722 executable mutators (319 for Rust, 403 for C++), each costing approximately \$0.08 using GPT-4o APIs, and targeting complex language constructs that are often overlooked by existing fuzzers.
  • We demonstrate the effectiveness of Mut4All by uncovering 96 compiler bugs (62 in rustc/gccrs, 34 in GCC/Clang), 54 of which have been confirmed as previously unknown, with 8 already fixed by developers. Additionally, Mut4All outperforms all baselines by finding more unique crashes and achieving the highest coverage on Rust and the second-highest on C++.
2. Illustrative Examples
To demonstrate Mut4All’s ability to generate sophisticated, feature-specific mutators, we present two confirmed bugs it uncovered in rustc and GCC.
2.1. An Example for Rust
Figure 1. Rust mutator specification generated by Mut4All referring to Rust-124182.
We first demonstrate how Mut4All generates mutators for Rust by showcasing a confirmed bug it uncovered in the official Rust compiler (rustc). The process begins by collecting historical bug reports from the Rust compiler’s official GitHub repository, each typically including a descriptive title, a triggering code snippet, compiler output, and a stack trace, as shown in the top half of Figure 1. These reports provide valuable insights into the compiler’s behavior and highlight language features that are insufficiently supported and may lead to new bugs.
Given this bug report, Mut4All invokes a pipeline of three LLM agents to generate a new mutator. The first agent, Mutator Invention, analyzes the bug description and the triggering code to identify key language features responsible for the bug. It then synthesizes a mutator specification comprising a detailed description, syntax correctness constraints, and a concrete Before and After example illustrating the transformation. As shown in Figure 1, the historical bug (Rust-124182333https://github.com/rust-lang/rust/issues/124182) is recognized by Mut4All as related to generic parameter handling and type layout, which is based on the presence of a generic type in the code snippet and a reference to type layout in the bug title. This insight guides the LLM agent to invent a mutator that duplicates a generic type and wraps it inside a tuple, as detailed in the specification in the bottom half of Figure 1. Note that this mutator is much more sophisticated and creative than those generated by existing methods such as MetaMut (Ou et al., 2024), which typically rely on a fixed set of syntactic transformations. In contrast, our approach leverages insights extracted from real-world bug reports to design mutators that go well beyond the expressive capabilities of these existing techniques.
1// Match generic type ‘T‘
2if let Type::Path(TypePath { path, .. }) = &first_field.ty {
3 if let Some(segment) = path.segments.first() {
4 if item_struct.generics.params.iter().any(|param| {
5 if let GenericParam::Type(type_param) = param {
6 type_param.ident == segment.ident
7 } else {
8 false
9 }
10 }) {
11 // Duplicate ‘T‘ and replace the type with ‘(T, T)‘
12 let generic_ident = segment.ident.clone();
13 first_field.ty = parse_quote!((#generic_ident, #generic_ident));
14 }
15 }
16}
Listing 1: The core implementation of the Rust mutator.
1trait MyTrait { fn virtualize(&self); }
2struct VirtualWrapper<T>(T, T); // From ‘(T)‘ to ‘(T, T)‘
3impl<T: static> MyTrait for T {
4 fn virtualize(&self) {
5 const { std::ptr::null::<VirtualWrapper<T>>() as *const dyn MyTrait };
6 }
7}
8fn main() { 0u8.virtualize(); }
Listing 2: The new bug Rust-14XX11 uncovered by Mut4All.
Next, the Mutator Implementation Synthesis agent, which is already enhanced via fine-tuning with a set of high-quality mutators, leverages the generated specification, our customized prompt templates, and several AST transformation examples extracted from Rust’s official AST library to produce a concrete mutator that performs the intended program transformation. The core logic of the resulting mutator is shown in Listing 1, which matches a generic type T at the AST level and replaces it with a tuple type (T, T). This transformation is designed to trigger compiler behavior related to generic types and type layout—the root cause of the original bug. In this case, the LLM generates a syntactically correct and semantically valid mutator without requiring any repair iterations. As a result, the Mutator Refinement agent is skipped for this example.
During our fuzzing phase, the mutator generates new program variants that stress the compiler’s type system in unexpected ways. As shown in Listing 2, Line 3 is mutated by introducing a self-referential tuple type VirtualWrapper<T>(T, T). This transformation causes the compiler to cast a pointer to the trait object dyn MyTrait into a pointer to the concrete type VirtualWrapper<T>, which is disallowed. The compiler fails to handle this edge case properly, resulting in an internal assertion failure. This bug, reported as Rust-14XX11, has been confirmed by Rust compiler developers as a previously unknown crash in rustc.
2.2. An Example for C++
Similarly, Mut4All generates mutators for C++ compilers by analyzing bug reports from LLVM’s GitHub repository. The process mirrors that of Rust, where LLM agents collaboratively invent, implement, and refine mutators based on compiler-specific insights. Figure 2 presents an example produced by Mut4All for C++.
The original bug report (LLVM-104802444https://github.com/llvm/llvm-project/issues/104802) describes a crash in the Clang C++ compiler triggered by a class that declares a member of its own type. In this case, the compiler fails to handle the invalid structure gracefully and enters infinite recursion during type checking. This suggests that recursive type patterns are not robustly guarded in certain compilation paths. Based on this insight, Mut4All’s Mutator Invention agent designs a mutator that intentionally injects self-referential type declarations, with the corresponding mutator specification shown in the bottom half of Figure 2. As mentioned above, existing methods such as MetaMut (Ou et al., 2024) are unable to design such mutators, as they rely on limited predefined syntactic transformations and do not account for language-specific features such as self-referential types.
Next, the Mutator Implementation Synthesis agent generates an implementation that inserts a self-referential field into a struct. Initially, the LLM produces a mutator using a deprecated AST API, resulting in a compiler error. After receiving the error feedback, Mut4All’s Mutator Refinement agent successfully repairs the mutator and produces a correct version (omitted here for brevity). The final mutator inserts a field such as S t; into a struct S, provoking recursive type behavior and potentially unsafe casting. Listing 3 shows an example of the mutated program, where the inserted self-reference interferes with a bit cast operation. This results in a compiler crash, reported as bug GCC-12XX98 which has been confirmed by GCC maintainers.
Figure 2. C++ mutator specification generated by Mut4All referring to LLVM-104802.
1struct S {
2 int s;
3 S t; // Inserted by Mut4All
4};
5auto t = __builtin_bit_cast (S, (0));
Listing 3: The new bug GCC-12XX98 uncovered by Mut4All.
3. Approach
Figure 3. The overview of Mut4All.
We now describe the design of Mut4All, with a detailed explanation of the processes underlying each step shown in the examples in Section 2. Figure 3 shows an overview of Mut4All, briefly describing the workflow and detailing the inputs and outputs of each component. Overall, Mut4All takes as input historical bug reports, prompt templates for each LLM agent, and existing manually written mutator examples, and produces a set of validated mutators ready for integration into our subsequent compiler fuzzing phase. At a finer granularity, as mentioned above, Mut4All comprises three specialized LLM agents corresponding to three key steps: (1) the Mutator Invention agent, which takes bug reports and bug-triggering code snippets as input and outputs a set of mutator specifications; (2) the Mutator Implementation Synthesis agent, which takes these specifications, along with our prompt templates and human-written mutator examples, to implement raw mutators in the target language; and (3) the Mutator Refinement agent, which takes the raw mutators and their associated compiler error messages (if any) as input and outputs validated mutators ready for large-scale compiler fuzzing. We now describe the design of each agent in detail.
3.1. Mutator Invention
Compilers, as the foundation of modern software development, are among the most actively maintained and scrutinized components in the open-source ecosystem. Bugs and usage issues are frequently reported on platforms such as GitHub, often accompanied by detailed discussions, triggering code snippets, and diagnostic outputs. These bug reports reflect valuable domain knowledge from compiler developers and often highlight specific language constructs or internal behaviors prone to failure. Mut4All leverages these historical bug reports as input to its first component, the Mutator Invention agent. This agent analyzes the bug descriptions and associated code to identify underlying language features responsible for the reported failures, and uses this analysis to design targeted mutation operators. It is based on the observation that many historical compiler bugs arise from recurring constructs or subtle corner cases that are difficult to expose via generic mutations—such as simple arithmetic rewrites or control-flow transformations. By grounding the mutation logic in real-world failure scenarios, Mut4All can generate semantics-aware, feature-specific mutators that are more likely to trigger previously unseen bugs.
To support this process, we design a structured prompt template (shown in Figure 4) that guides the LLM to extract relevant language features from each bug report and synthesize a precise, context-aware mutation operator. The prompt is designed with clear functional instructions: it explicitly requires the LLM to avoid superficial edits, target well-defined code constructs, and design mutations that achieve high coverage across diverse code contexts. To improve generalizability, it discourages the use of ad hoc identifiers and promotes the reuse of existing code elements.
[Role]: You are a {Language} compiler expert. [Task]: Here is a bug report for the {Language} compiler labeled by {Tag}. Summarize the structures that [Language} compiler bugs are more likely to trigger, and design mutation operators… [Functional Instructions]: 1. Target a specific code construct and clearly state where it applies; 2. Customized data names, types, and variable names…; 3. Achieve high coverage across different program structures and features; 4. … [Output Instructions]: The output must be plain text. Do not use the Markdown format. Design exactly one mutation operator per request… [Input]: Here are the contents of the bug report: {Bug Descriptions} and {Bug-Triggering Code}.
Figure 4. Prompt template for mutator invention.
3.2. Mutator Implementation Synthesis
Once a mutator has been invented, the next step is to synthesize its executable implementation. This typically involves writing AST transformation code that modifies a program according to the intended mutation—a process that is tedious and error-prone when done manually, especially for complex and evolving compiler infrastructures. To address this, Mut4All introduces a second agent responsible for Mutator Implementation Synthesis. This LLM-based agent generates transformation code tailored to the target language based on the mutator specification. To ensure code quality and semantic correctness, we fine-tune the LLM using OpenAI’s standard fine-tuning API, guided by 10 human-written mutator examples per language. These examples cover diverse syntactic constructs and representative transformation patterns, serving as domain-specific demonstrations that help the model synthesize accurate and idiomatic AST manipulation code.
To enforce structural consistency, Mut4All adopts a predefined mutator template for each supported language. The template defines the skeleton that the LLM must complete with appropriate mutation logic like matching AST nodes, applying transformations, and handling edge cases. Listing 4 and Listing 5 present simplified examples for Rust and C++, respectively. Each template specifies the mutator’s entry point, filtering conditions, and placeholders for inserting LLM-generated transformation code.
As illustrated in Figure 3, the Mutator Implementation Synthesis agent combines three key ingredients: (1) the mutator specification, (2) the language-specific mutator template, and (3) human-crafted AST transformation examples. It invokes the fine-tuned LLM with a structured prompt (shown in Figure 5) that clearly defines the agent’s role, functional constraints (e.g., correct API usage, no syntax errors), and output format. The output of this agent is a raw mutator implementation, which is subsequently passed to the Mutator Refinement agent for potential repair and validation.
[Role]: You are an expert in {Language} compiler development and AST manipulation. [Task]: Based on the provided mutator specification and example mutators for the target language, complete the {Language} mutator template to implement the intended transformation. [Functional Instructions]: 1. Ensure that the generated code is syntactically valid {Language}; 2. Correctly implement the mutation logic as described in the specification; 3. Follow the structure of the provided template; 4. Use appropriate {Language} APIs for AST manipulation; 5. Includes all necessary imports and dependencies… [Output Instructions]: Output only the completed {Language} code, without any additional explanations or formatting… [Input]: Here are the necessary inputs: Mutator Specification: {Mutator Specification}. Mutator Template for {Language} : {Mutator Template}. Human-written {Language} mutators: {Mutator Examples}.
Figure 5. Prompt template for mutator implementation synthesis.
1impl Mutator for Mutator_ {
2 fn name(&self) -> &str {
3 "Mutator_"
4 }
5 fn mutate(&self, file: &mut syn::File) {
6 for item in &mut file.items {
7
8 }
9 }
10 fn chain_of_thought(&self) -> &str {
11 ""
12 }
13}
Listing 4: The core part of the mutator template for Rust.
[…]
2 if (auto *MT = Result.Nodes.getNodeAs<clang::{{ASTNodeType}}>({{NodeName}})) {
3 if (!MT || !Result.Context->getSourceManager().isWrittenInMainFile(MT->getLocation()))
4 return;
5 }
6}
Listing 5: The core part of the mutator template for C++.
3.3. Mutator Refinement
In this agent, each raw mutator generated in the above stage undergoes validation and potential repair by the Mutator Refinement agent. Mut4All begins by extracting all bug-triggering programs from the original bug reports and assembling them into a seed test suite. These programs are used to validate whether the mutator (1) compiles successfully, (2) integrates with the compiler’s AST infrastructure, and (3) performs its transformation without introducing compilation failures in downstream programs.
If one mutator fails to compile, or if applying it to any seed program causes a compilation error, Mut4All captures the full compiler error messages (e.g., due to missing imports, incorrect type handling, or invalid AST rewrites). These messages are embedded into a structured prompt template (Figure 6) that instructs the LLM to revise the mutator while preserving its intended transformation semantics. The refinement process is conducted in an iterative loop. After each repair attempt, the revised mutator is recompiled and re-applied across all seed programs. If the mutator passes compilation and does not cause any errors when applied to the test suite, it is marked as valid and added to the mutator pool for fuzzing. Otherwise, the process continues for up to a maximum of $N$ iterations (we use $N=10$ in our experiments). If the mutator still fails validation after the limit is reached, it is discarded.
Furthermore, to ensure that the mutator performs meaningful transformations, we include a post-validation check: if the mutator compiles and runs but does not introduce any syntactic change in any test input (e.g., due to overly strict matching logic), it is considered ineffective and also discarded. This ensures that the final mutator set is not only valid, but actively contributes to provide meaningful and diverse mutations for the fuzzing phase.
[Role]: You are a {Language} compiler expert revising a mutator code to fix errors while preserving its mutation logic. [Task]: Given the buggy mutator code and the associated compiler error messages, produce a corrected version that resolves the errors without changing the core mutation behavior. [Functional Instructions]: 1. Prioritize fixing the provided compiler errors by identifying their causes (e.g., missing imports, type mismatches) and applying targeted corrections. 2. Ensure the revised code is syntactically valid {Language} and compiles successfully. 3. … [Output Instructions]: Output only the corrected {Language} code in plain text, without any explanations, comments, or Markdown formatting. [Input]: Here are the necessary inputs: Previous Mutator Code: {Previous Code} Compile Errors: {Compile Errors}
Figure 6. Prompt template for mutator refinement.
3.4. Fuzzing Loop
After collecting a set of validated mutators, Mut4All enters its core fuzzing loop to systematically test target compilers. Algorithm 1 outlines this process in detail.
Given an initial set of seed programs $S$, Mut4All first expands the seed corpus using rule-based transformations via the enhance function, following the fixed-replacement strategy introduced by Clozemaster (Gao et al., 2025), which replaces each identifier in the seed programs with a randomly chosen identifier of the same type. This step aims to diversify the seed space and uncover additional transformation opportunities. The enhanced seed corpus is denoted as $\mathcal{P}$. In each fuzzing iteration, Mut4All randomly selects $i$ mutators from the validated mutator set $M$ and applies them to each program in the current seed pool $\mathcal{P}$. The resulting mutated programs $p^{\prime}$ are then compiled and executed on each target compiler $c$ in the compiler set $C$. If a compiler crashes or hangs, the corresponding input is immediately recorded as a bug-triggering example. For successfully compiled programs, the execution results $r$ of the generated binaries are collected. Mut4All compares these outputs across compilers or against expected values (when available). If any inconsistency is detected—such as divergent outputs, abnormal termination, or unexpected side effects—a behavioral bug is reported. The fuzzing loop continues until a predefined timeout $T$ is reached or until other stopping criteria (e.g., maximum number of iterations or detected bugs) are met.
1
Input: Seed Programs: $S$
Input: Valid Mutators: $M$
Input: Compilers Under Test for a Certain Language: $C$
Input: Timeout: $T$
2
[…]
4 while $\operatorname{time\_not\_expired}$ ($T$) do
5  for $i\leftarrow 1$ to $\operatorname{Int.Max}$ do
6     $M^{\prime}\leftarrow\operatorname{random\_select\_N\_mutators}$ ($M$, $i$)
7     foreach $p\in\mathcal{P}$ do
8        foreach $m\in M^{\prime}$ do
9           $p^{\prime}\leftarrow\operatorname{apply}(p,m)$
10
11          $R\leftarrow\phi$
12           foreach $c\in C$ do
13              $\operatorname{res,exe}\leftarrow\operatorname{compile}$($c$, $p$)
14              if $\operatorname{crash(res)}~OR~\operatorname{hang(res)}$  then
[…]
16
17             else if $\{\operatorname{exe}\}\neq\phi$ then
18                 $r\leftarrow\operatorname{execute}$($\operatorname{exe}$)
19                 $R\leftarrow R\cup\{\operatorname{res}\}$
20
21
22          if $\operatorname{has\_inconsistency}$ ($R$) then
23              $\operatorname{record}$ ($p$, $R$)
24
25
26
27
28
Algorithm 1 The fuzzing loop of Mut4All.
4. Evaluation
Our evaluation aims to answer the following research questions:
  1. RQ1
    How does Mut4All compare to the state-of-the-art compiler fuzzers?
  2. RQ2
    Is Mut4All effective in finding real-world bugs in production-level compilers?
  3. RQ3
    How effective does Mut4All in mutator synthesis for compiler fuzzing?
4.1. Implementation
To demonstrate the cross-language capability of Mut4All, we select two prominent system-level languages, Rust and C++, as a proof of concept. We implement a dedicated web crawler to extract fixed or resolved bug reports from the official GitHub repositories [^5] https://github.com/rust-lang/rust and https://github.com/llvm/llvm-project of rustc and Clang, and parse their contents to extract the embedded bug-triggering programs. For each target language, we construct the initial seed corpus by combining all extracted bug-triggering programs with the official test suites of the corresponding compilers (Zhong, 2022). This process yields 20,527 seeds for Rust and 27,786 seeds for C++. We expand the corpus to 100,000 programs using Clozemaster-style rule-based mutations (Gao et al., 2025), which replace each identifier in the seed programs with a randomly selected identifier of the same type to diversify the seed space. We then discard programs that fail to compile, resulting in 87,733 Rust and 68,176 C++ programs. Additionally, from the crawled bug reports, we randomly select 500 reports per language for the Mutator Invention agent to synthesize language-specific mutators. In total, Mut4All generates 319 valid mutators for Rust and 403 valid mutators for C++.
We adopt GPT-4o (Achiam et al., 2023) as the underlying LLM, which also offers an API for custom fine-tuning. We utilize this capability to further improve the quality and correctness of the generated mutators by providing 10 manually crafted examples that cover a diverse range of AST transformations as additional fine-tuning data for each language.
For the bug-hunting experiments, we allocate a fuzzing time budget of 48 hours per language and initiate the fuzzing loop using all valid mutators. Within this time budget, we observe that the fuzzing pipelines for both Rust and C++ typically remain in the stage of applying up to two mutators per seed.
4.2. Evaluated Languages and Compilers
Table 1. Evaluated languages, compilers, and baseline tools.
Language Compiler Fuzzer Avenue
Rust rustc/gccrs RustSmith (Sharma et al., 2023) ISSTA-23
Rustlantis (Wang and Jung, 2024) OOPSLA-24
Clozemaster (Gao et al., 2025) ICSE-25
C++ GCC/Clang CSmith (Yang et al., 2011) PLDI-11
YARPGen-2.0 (Livinskii et al., 2023) PLDI-23
GrayC (Even-Mendoza et al., 2023) ISSTA-23
Fuzz4All (Xia et al., 2024) ICSE-24
MetaMut (Ou et al., 2024) ASPLOS-25
For Rust, we adopt rustc [^6] https://www.rust-lang.org/ (the official compiler) and gccrs [^7] https://github.com/Rust-GCC/gccrs/ (the GNU front-end for Rust). For C++, we use GCC [^8] https://gcc.gnu.org/ and Clang, [^9] https://clang.org/ which are the two most widely deployed C++ compilers in practice. We use the latest release versions of all compilers: rustc 1.88, gccrs-a1a56c6, GCC 14.1.0, and Clang 18.1.0. As test oracles, we apply crash, hang, and differential testing for compilers of the same language (Gao et al., 2025; Zhang et al., 2017). Among them, crash and hang are used to detect compiler failures, the former indicates a crash or assertion failure, while the latter indicates an infinite loop or deadlock. Differential testing is used to detect inconsistent behaviors between different compilers, which is a common practice for testing C/C++ compilers that share the same language standard but may have different implementations. For the baseline fuzzers, we use their default configurations from their official repositories. For approaches that require seed programs, we use the same initial seed corpus as Mut4All to ensure a fair comparison (i.e., the enhanced seeds).
4.2.1. Baselines
To thoroughly evaluate the effectiveness of Mut4All, we compare it against both traditional compiler fuzzers and recent LLM-based approaches. In particular, to investigate how our mutator-driven pipeline compares to direct LLM-based mutated program generation, we include the latest LLM-based fuzzers that modify program source code directly—namely, Clozemaster (Gao et al., 2025) for Rust and Fuzz4All (Xia et al., 2024) for C++. In total, we evaluate three fuzzers for Rust and five for C++, as shown in Table 1.
Traditional Fuzzers.  For Rust, we adopt generation-based fuzzers such as RustSmith (Sharma et al., 2023) and Rustlantis (Wang and Jung, 2024). For C++, we include CSmith (Yang et al., 2011) and YARPGen (Livinskii et al., 2020, 2023), which follow a similar generation-based paradigm by randomly expanding language grammars and production rules to produce valid programs. Additionally, we include GrayC (Even-Mendoza et al., 2023), a representative mutation-based C++ fuzzer equipped with manually designed, general-purpose mutators.
LLM-based Fuzzers.  We further compare Mut4All with recent LLM-based compiler fuzzers that follow different design philosophies: we let the LLMs design and implement mutators, which are then applied to the seed programs but they let the LLMs directly edit the source code to generate mutated programs. For Rust, Clozemaster (Gao et al., 2025) fine-tunes an LLM to complete masked code segments, enabling direct in-place edits to source programs. For C++, Fuzz4All (Xia et al., 2024) utilizes LLMs to rewrite seed programs guided by language standards and official API references. In contrast, MetaMut (Ou et al., 2024) uses an LLM to systematically synthesize and combine simple transformation rules into new mutators, falling into the same category as Mut4All but with limited mutator diversity.
4.3. RQ1: Comparison with Existing Fuzzers
To evaluate Mut4All against existing approaches, we run each fuzzer with a 24-hour time budget, which is a common setting in prior work (Even-Mendoza et al., 2023; Xia et al., 2024; Ou et al., 2024; Yang et al., 2024a). Also, each fuzzer is executed independently for five trials. Fuzz4All and MetaMut both use GPT-4o as Mut4All, while Clozemaster employs a fine-tuned model. We execute each fuzzer five times independently.
4.3.1. Unique Crashes
Following MetaMut (Ou et al., 2024), we identify a unique crash based on the top-2 frames of the stack trace, excluding compiler helper functions. We aggregate all unique crashes across the five runs for each fuzzer, as shown in Figure 7.
Mut4All consistently discovers the highest number of unique crashes for both Rust and C++, significantly surpassing all baselines. For Rust (Figure 7(a)), Mut4All uncovers 60 unique crashes, whereas Clozemaster and Rustlantis report only 5 and 0, respectively. Notably, only 6 crashes overlap between Mut4All and Clozemaster, indicating that the vast majority of bugs found by Mut4All are exclusive. For C++, Mut4All detects 35 unique crashes, compared to 27 from MetaMut and 10 from Fuzz4All. The overlap is minimal—only 3 crashes are shared between Mut4All and each of these fuzzers—demonstrating that most of Mut4All’s findings are novel. In contrast, traditional fuzzers such as CSmith, GrayC, and YARPGen report no unique crashes in our experiments, suggesting that these conventional techniques are reaching their saturation point.
(a) Rust
(b) C++
Figure 7. Venn diagram of discovered unique crashes.
4.3.2. Coverage
(a) rustc
(b) GCC
(c) Clang
Figure 8. Line coverage over 24h of fuzzing.
Figure 8 presents the average coverage trends of all evaluated fuzzers. For Rust (Figure 8(a)), Mut4All achieves the highest line coverage on rustc, outperforming Clozemaster and Rustlantis. For C++ compilers, MetaMut achieves the highest code coverage on both GCC (Figure 8(b)) and Clang (Figure 8(c)). Mut4All ranks second, followed closely by Fuzz4All. This result aligns with expectations, as MetaMut employs a coverage-guided fuzzing strategy that explicitly optimizes for maximizing code paths executed. In contrast, Mut4All focuses on synthesizing and evaluating a diverse set of mutators across seed programs, without using coverage feedback as a direct guidance signal. Despite this, Mut4All still achieves competitive coverage when compared to state-of-the-art fuzzers, particularly in Rust. Mut4All ’s diverse mutators improve coverage without explicit optimization for coverage.
Answer to RQ1: Mut4All consistently outperforms all baselines in finding unique crashes, and achieves the best coverage on Rust and the second-best on C++.
4.4. RQ2: Bug-Hunting Effectiveness
4.4.1. The Status of Our Reported Bugs
We have synthesized 319 and 403 valid mutators for Rust and C++, respectively. Using these mutators, we conducted a 48-hour fuzzing campaign for each language and reported a total of 96 bugs, as summarized in Table 4.4.1. Specifically, we identified 62 bugs in Rust compilers (44 in rustc and 18 in gccrs) and 34 in C++ compilers (21 in GCC and 13 in Clang). Among these, 54 bugs (56%) have been confirmed by compiler maintainers as previously unknown bugs: 33 in rustc, 5 in gccrs, 6 in GCC, and 10 in Clang. Notably, 8 of the confirmed bugs have already been fixed by the respective compiler developers. In addition, 14 of our reports were classified as duplicates of existing bugs, and 2 rustc reports were marked as Won’t Fix.
These results suggest that Mut4All can uncover a significant number of bugs within a limited time budget and while applying at most two mutators per seed program, demonstrating the efficiency and effectiveness of our mutator-driven approach.
Table 2. Status of the reported bugs by Mut4All
Language Rust C++ Total
Compiler rustc gccrc GCC Clang
Reported 44 18 21 13 96
Confirmed as New 33 5 6 10 54
Fixed 4 3 0 1 8
Duplicate 9 0 2 3 14
Won’t Fix 2 0 0 0 2
Invalid 0 0 0 0 0
Unconfirmed 0 13 13 0 26
4.4.2. Bug Examples
We further present two more bugs to show Mut4All’s effectiveness in uncovering subtle, hard-to-construct compiler issues beyond prior fuzzers.
1trait Trait<V, D> {}
2struct A;
3struct B;
4impl<V, D> Trait<X, X> for A // ‘Trait<V, D>‘ -> ‘Trait<X, X>‘
5where A: Trait<D, V>, B: Trait<V, D>, {}
6impl<V, D> Trait<X, X> for B // ‘Trait<V, D>‘ -> ‘Trait<X, X>‘
7where A: Trait<V, D> {}
8fn impls_trait<T: Trait<V, D>, V, D>() {}
9fn main() { impls_trait::<B, _, _>(); }
Listing 6: The new bug Rust-14XX63 uncovered by Mut4All.
Listing 6 shows a Rust program where valid trait parameters are replaced with an undefined type (i.e., X), causing rustc to enter an infinite loop due to unresolved recursive bounds. While such mutations are unlikely to be crafted manually, Mut4All automatically identifies them, exposing a deep bug in the compiler’s trait resolution logic.
1template<typename T>
2// ‘fooT()‘ -> ‘fooT(C auto,)‘
3constexpr int fooT(C auto,) { return __alignof__(T); }
4template <typename T>
5struct __attribute__((aligned(fooT<T>()))) S { char dummy; };
6S<int> s;
Listing 7: The new bug LLVM-14XX70 uncovered by Mut4All.
Listing 7 presents a C++ program using template and constexpr. A mutator from Mut4All modifies a valid function fooT() into one with an invalid parameter list: fooT(C auto,). Though syntactically incorrect, this triggers a Clang crash, exposing a long-standing bug dating back to version 9.0 (i.e., more than 6 years). Such unintuitive mutations in sensitive constructs are rarely produced by traditional fuzzers like MetaMut, nor by human developers. Mut4All’s systematic exploration of these edge cases demonstrates its ability to uncover deep compiler bugs.
Answer to RQ2: Mut4All discovered 96 bugs across four production compilers, 54 of which have been confirmed as previously unknown, demonstrating its effectiveness in hunting real compiler bugs.
4.5. RQ3: Mutator Synthesis Effectiveness
The mutators designed and implemented in Mut4All have shown their effectiveness in fuzzing compilers. In this RQ, we evaluate the overall generation cost and the common problems in valid/invalid mutators, which can help us understand the effectiveness of Mut4All in mutator synthesis.
4.5.1. Overall Effectiveness and Cost
We sampled 500 historical bug reports for both Rust and C++ compilers, and used them to construct prompts for Mut4All to invent and synthesize mutators. We then performed a statistical analysis of the synthesis outcomes.
For Rust, Mut4All generated 64% valid mutators (319 out of 500). Among invalid ones, 110 were filtered out for producing no observable mutations on the test suite, and 71 failed to compile after 10 rounds of refinement. Notably, 397 of the Rust mutators required at least one round of refinement, with an average of 3.51 refinement iterations per mutator. On average, generating one valid Rust mutator consumed 14,969 input tokens and 3,694 output tokens, corresponding to \$0.074 each based on the GPT-4o pricing in 2025.
For C++, Mut4All synthesized 403 valid mutators (81%), while 9 were ineffective and 88 failed after 10 refinement attempts. Among these, 264 (52.8%) required refinement, with an average of 2.48 iterations. The generation of each valid C++ mutator consumed an average of 22,276 input tokens and 2,500 output tokens, amounting to a cost of approximately \$0.080.
Overall, Mut4All achieved a higher success rate on C++ than on Rust, primarily due to C++’s more concise and high-level AST transformation APIs. In contrast, Rust’s verbose, low-level visitor-based traversal increases the likelihood of failure. Importantly, since the entire synthesis process is fully automated and requires no human intervention, the cost of approximately \$0.08 per valid mutator is highly affordable and practical for large-scale use.
4.5.2. Analysis of Valid Mutators
We manually analyzed 65 Rust and 68 C++ mutators, a statistically representative number with a 90% confidence level and a 10% error margin. Two authors manually review each mutator and examined the alignment between code logic and the corresponding specifications.
The analysis showed that only 35% of the Rust mutators and 28% of the C++ mutators faithfully implemented the intended mutations. Interestingly, among the mutators that triggered confirmed bug discoveries, approximately 65% were fully aligned with their specifications. This suggests that even partially incorrect mutators can uncover bugs! Nonetheless, higher implementation fidelity is generally associated with improved bug-finding ability, underscoring the importance of enhancing mutator correctness in future work.
To further understand the the errors in the mutators, which guide the direction to improve our method, we categorized the implementation errors into three types: (1) Missing changes (19%): the specification involves changes at multiple locations, but some were omitted; (2) Incorrect mutation (45%): the code transformation was implemented incorrectly; (3) Incorrect matching (36%): the mutator failed to match the correct AST nodes.
4.5.3. Analysis of Invalid Mutators
Using the same sampling parameters (90% confidence level, 10% error margin), we analyzed 45 Rust and 43 C++ invalid mutators. Via open coding, we identified 5 primary causes of failure: (1) Use of deprecated APIs (40%), (2) Incorrect parameters (39%), (3) Syntax parsing errors (10%), (4) Operand type mismatch (7%), and (5) Use of undeclared names (5%).
The most common failure in Rust was Incorrect parameters (62%), while in C++ it was the Use of deprecated APIs (65%). These results indicate that synthesizing structurally correct and type-safe AST manipulations remains challenging for advanced LLMs, particularly in languages with verbose or rapidly evolving APIs like Rust.
Answer to RQ3: Mut4All automatically synthesizes valid and effective mutators with low cost (about \$0.08 each) without human intervention. Improving the correctness of LLM-synthesized mutators remains promising for future work.
5. Discussion
5.1. Limitation and Implication
Our study demonstrates the effectiveness of leveraging LLMs to invent and implement compiler fuzzing mutators. With carefully designed prompts and systematic validation mechanisms, Mut4All functions as a fully automated, end-to-end framework, offering a compelling alternative or complement to manually crafted mutators. It can generate a large number of mutators at relatively low cost, significantly reducing human effort and enabling scalable mutation design.
Despite these advantages, our evaluation also reveals several key limitations. First, only about 30% of the valid mutators fully adhere to their intended specifications. Among those that triggered confirmed bugs, about 65% were correctly implemented. This suggests that while imperfect mutators may still uncover bugs by chance, higher implementation accuracy generally correlates with better bug-finding effectiveness. Improving the semantic fidelity of generated mutators remains a critical direction for future work. Moreover, many invalid mutators fail due to deprecated APIs, likely inherited from legacy training data. To enhance correctness, future work could explore symbolic reasoning or reinforcement learning and fine-tune models on modern code.
5.2. Threats to Validity
Internal Validity.  Our pipeline depends on LLM output quality and prompt design. Although we employ validation and refinement to ensure correctness, LLM-generated mutators may still misbehave in edge cases. Manual analysis may also introduce bias, which we mitigate using statistically sound sampling (90% confidence, 10% margin of error).
External Validity.  We evaluate on Rust and C++, using four widely adopted compilers. While representative, results may not generalize to other languages or domains. Baselines span both traditional and LLM-based fuzzers, but future tools may differ in design.
Construct Validity.  We primarily use crash count and line coverage as evaluation metrics. While standard in compiler fuzzing, these do not capture semantic correctness or non-crashing logic bugs. Additionally, fuzzers are executed under timeout constraints, which may introduce variance; we mitigate this by averaging results across multiple independent runs, i.e., 5 runs for each fuzzer.
6. Related Work
6.1. Compiler Fuzzers
Compiler fuzzers have uncovered numerous bugs in production-level compilers (Chen et al., 2020; Xu et al., 2023; Marcozzi et al., 2019; Chen et al., 2016a), many of which are critical (Marcozzi et al., 2019; Zhong, 2025) and can adversely impact downstream software systems. Broadly, compiler fuzzing techniques fall into two main categories: generation-based and mutation-based approaches.
Generation-based fuzzers synthesize programs from scratch using grammar rules or probabilistic models. Tools such as CSmith (Yang et al., 2011) and YARPGen (Livinskii et al., 2020, 2023) exemplify this paradigm by generating semantically valid C/C++ code. Recent extensions have expanded to other languages, including Java (Chen et al., 2023; Zhao et al., 2022), JavaScript (Ye et al., 2023), Rust (Sharma et al., 2023; Wang and Jung, 2024), MLIR (Wang et al., 2023a), deep learning compilers (Wang et al., 2023b; Liu et al., 2023a; Ren et al., 2025), SQL (Hua et al., 2023), and Kotlin (Georgescu et al., 2024). LLM-based generators (Yang et al., 2024a; Gao et al., 2025) further automate program synthesis but struggle with intricate language constructs (e.g., generics, trait bounds), which are difficult to generate from scratch while preserving semantic correctness. In contrast, Mut4All mutates real-world programs—retaining structural complexity and ensuring semantic validity—thereby enabling deeper and more effective exploration of compiler internals.
Mutation-based fuzzers generate program variants by applying transformations to seed programs. Representative approaches include EMI-based fuzzers (Le et al., 2014, 2015; Gao et al., 2024b) and differential fuzzers targeting C/C++ (Even-Mendoza et al., 2023; Jiang et al., 2021; Liu et al., 2023b), JIT languages (Chen et al., 2016b, 2019; Schumi and Sun, 2021; Wu et al., 2023; Park et al., 2023; Gao et al., 2024a), deep learning compilers (Li et al., 2023; Chen et al., 2025b), MLIR (Suo et al., 2024), Rust (Yang et al., 2024b), and various domain-specific languages (Tu et al., 2022; Bendrissou et al., 2024). These methods typically rely on manually designed mutators, requiring significant domain expertise and engineering effort. Among them, MetaMut (Ou et al., 2024) represents the first attempt to synthesize mutators using LLMs, composing transformations from predefined AST-level operations. However, its search space is constrained by these predefined operations, and over half of the generated mutators require manual repair to function correctly. In contrast, Mut4All automates both the invention and implementation of mutators through a multi-agent LLM framework guided by real-world bug reports crawled from open-source compiler communities. By grounding mutation design and synthesis in authentic bug-prone features and enabling collaboration among specialized LLM agents, Mut4All explores a broader mutator design space with greater creativity and incorporates automatic validation and refinement mechanisms to ensure correctness, significantly reducing human effort while improving performance.
6.2. LLM-Based Testing
Recent advances have demonstrated the potential of LLMs in automated software testing (Wang et al., 2024; Chen et al., 2025a). Many works focus on generating unit tests, including test functions (Schäfer et al., 2023; Chen et al., 2024a; Yang et al., 2024c; Yuan et al., 2024; Shang et al., 2025), input values (Chen et al., 2024b), and assertion oracles (Hossain et al., 2025; Molina et al., 2025), often guided by code structure, comments, or coverage feedback. Bug reports are used to enhance test generation (Ouédraogo et al., 2025). These approaches primarily operate at the unit-test level, aiming to validate method-level behavior, and their outputs are usually test code written in the same language as the program under test. In contrast, our work targets compiler testing, focusing on mutation operators that systematically mutate code to explore deep compiler behavior.
7. Conclusion
We present Mut4All, a fully automated, language-agnostic framework that synthesizes mutation operators using LLMs guided by historical bug reports. Mut4All uses three LLM agents for mutator invention, implementation, and refinement. It generates 319 and 403 valid mutators for Rust and C++, respectively, which uncovered 96 compiler bugs, 54 of which were previously unknown. The low cost demonstrates the promise of LLM-enhanced pipelines as a practical alternative to manual mutator development. In the future, we plan to reduce invalid mutators and enhance the correctness of mutator implementation by symbolic reasoning or reinforcement learning.